# Stable self-adaptive timestepping for Reduced Order Models for incompressible flows

### v1

@@ 4 Numerical experiments
In this section, the new adaptive timestepping method outlined in Algorithm 1 is tested for two different fluid flow test cases. First of all, the robustness of the method in a case without boundary conditions, namely the shear-layer roll-up, is presented. Moreover, the evolution of the maximum stable $\Delta t$ for different numbers of modes is presented, and compared with the maximum stable $\Delta t$ using the AlgEigCD. In the second case, the influence of non-homogeneous boundary conditions on the preservation of stability is considered for the unsteady wake behind an actuator disk [8]. In both test cases, the classical explicit fourth-order RK scheme is used in a two-dimensional incompressible Navier-Stokes solver [52, 53]. Both methods have been run for $M=[16,32,64,128,200]$ modes. Table 2 summarizes the methods used in this work.
@@
C ROM+constant ROM with constant $\Delta t$; used to compare accuracy of ROM+RedEigCD.
D Best ROM Best approximation ROM (Eq. (39)) with constant $\Delta t$;
used to verify $\Delta t_{\text{ROM}}>\Delta t_{\text{FOM}}$ throughout simulation.
@@ 4.1 Shear-layer roll-up
Figure 5 (left) shows that the new self-adaptive timestepping method is effective in providing a larger timestep for the ROM than for the FOM, as expected from Theorem 1, as even for $M=200$, the largest number of modes considered, the ratio is around 1, which is indicated in the plot with a horizontal thicker line. Depending on the current flow physics and number of modes of the ROM, this timestep can be up to 9 times larger than for the FOM, which significantly accelerates the simulation by notably reducing the number of timesteps required to reach the final time. Moreover, as shown in Figure 7, the accuracy of the ROM is not compromised by the use of RedEigCD.
Furthermore, Figure 5 (left) shows how the ratio between the timesteps obtained in Case B and Case A converges to 1 for increasing ROM dimension, as expected from Theorem 1, as the eigenvalues of the reduced operators interlace with those of the full-order operators. This is especially lucid for $M=200$, where the singular values have decayed to $\mathcal{O}(10^{-9})$, and the truncation errors between the ROM and the FOM (Figure 4) get closer to zero. Furthermore, note that at certain points, especially at the beginning of the simulation, the timestep ratio for $M=200$ is slightly smaller than 1.
@@
Figure 7 shows that the errors for both Case B and Case C have the same magnitude regardless of the number of modes ($M=16$, left; $M=200$, right). Thus, the use of the self-adaptive timestepping methods as performed in these cases does not significantly affect the accuracy of the ROM compared to the error of the best approximation ROM, as shown in Figure 7. In this case, $\Delta t$ for Case C is set to 0.01, as in the test case from [8]. The used error is defined as follows,
@@
where $\mathbf{u}_{\mathrm{test}}$ indicates the velocity field used to compute the error, i.e. $\mathbf{u}_{\mathrm{\texttt{B}}}$, $\mathbf{u}_{\mathrm{\texttt{C}}}$ and $\mathbf{u}_{\mathrm{\texttt{D}}}$; $\mathbf{u}_{\mathrm{\texttt{A}}}$ corresponds to the snapshot at that given point in time, and $||.||_{X}$ corresponds to the $X$-weighted $L_{2}$ norm.
Figure 7: Error comparison of the ROM results for $M=16$ (left) and $M=200$ (right) in the shear-layer roll-up of $\text{Re}=1000$ for Case B, C, and D.
In terms of efficiency, the ROM with $M=64$ reduced the number of unknowns from $3\times 10^{4}$ to 64, while simultaneously allowing timesteps up to almost 5 times larger than the FOM. This reduced the total number of timesteps, while the wall-clock time for each timestep was also reduced due to dimensionality reduction, thus having benefits from both factors. Figure 7 (left) shows that for $M=16$ there is no loss of accuracy for the new self-adaptive timestepper (Case B) compared to Case C, which indicates that in this test case there is no loss of accuracy regardless of the number of modes, yet for this case the benefits of RedEigCD are greater, with potential gains of up to a factor of 9.
Moreover, the accuracy of estimating the eigenvalues for the convective term using RedEigCD has been tested against the actual eigenvalues. To do so, the reduced convective operator for the FOM at $t=20$ has been constructed and its eigenbound has been precisely calculated and compared to the one estimated using RedEigCD, as reported in Section 4.3.
@@ 4.2 Flow through an actuator disk
In this case, the air flow around a wind turbine is modelled as in [22]. The simulation geometrical domain is $0,10]\times[-2,2]$, with a time-dependent prescribed inlet velocity at the left boundary, as defined as follows,
@@
The other boundaries correspond to an outflow condition [52, Section 6.4.2]. The wind turbine has been modelled using an actuator disk of diameter 1, placed at $(2,0)$, which corresponds to a momentum sink modelled as a constant force $f=0.25$ acting in the negative $x$-direction. The initial conditions for this case are
@@
The magnitude of the ROM error, calculated with Eq.(40), is not affected by the variable $\Delta t$ compared to the reference result from [8], where $\Delta t$ was set to $4\pi/200$. As shown in Figure 10, the accuracy of the ROM+RedEigCD (Case B) is comparable to that of the ROM with constant $\Delta t$ (Case C), for both $M=16$ and $M=200$. Note that the error for Case D (best approximation ROM) is significantly smaller than for Cases B and Case C, as expected.
Figure 10: Error comparison of the ROM results for $M=16$ (left) and $M=200$ (right) in an actuator disk simulation with $\text{Re}=100$ for Case B, Case C, and Case D.
@@ 4.3 Accuracy of the eigenbound estimates
In order to compute an overall error between the actual eigenvalues and the estimations, the following error is defined:
@@ 5 Conclusion
In this paper, the application of self-adaptive timestepping for the time integration of a POD-Galerkin ROM for the incompressible Navier-Stokes simulations has been presented. The method is based on the conceptual ideas from EigenCD [6] and AlgEigCD [7], in which the eigenbound of the convective and diffusive operators is estimated in such a way that the eigenbound of the convection-diffusion operator can be placed on the boundary of the stability region of the explicit time integration scheme used.
Nonetheless, when the aforementioned methods would be applied directly to a ROM framework, they would require operations that scale with the FOM dimension. To avoid these operations, we have proposed a new methodology, RedEigCD, that can effectively estimate the eigenbound of reduced convective and diffusive operators. We extend upon the methodology applied by Sanderse [8], where the reduced convective operator is split into a quadratic and a linear contribution with the ROM coefficients, and utilized the sub-additivity of eigenbounds as well as Bendixson’s rectangle. Thus, this allowed simplifying the method to a linear combination of the eigenbounds of the reduced convective operator, which is computed offline. On the other hand, the diffusive operator does not change in time. Therefore, its eigenbound can be computed exactly in an offline phase as well. The combination of these two computations, along with scaling the timestep as in [6, 7] with this eigenbound, forms the core of RedEigCD.
The presence of non-homogeneous boundary conditions that affect the linear term in the reduced convective term is also considered. Their interactions are modifying the structure of the matrix, breaking its skew-symmetry arising from the symmetry-preserving ROM that was applied. We proposed to address this by splitting this contribution into its symmetric and skew-symmetric parts, and adding them to the diffusive and convective terms, respectively.
The method has been tested in a fully-periodic shear-layer roll-up for up to $t=20$, which tests the general aspects of the methodology while avoiding any boundary conditions. It has been shown that, in general, the $\Delta t$ obtained for the ROM using RedEigCD is larger than the $\Delta t$ at the FOM level, due to the interlacing of eigenvalues, which leads to smaller eigenbounds for the reduced counterparts. The benefits in timestep size for the ROM+RedEigCD for the shear-layer roll-up can be close to a factor of 10, depending on the number of modes used and the flow conditions. This leads to reduced simulation times due to the decreased number of timesteps needed.
Moreover, the method has been tested in a case with time-dependent boundary conditions. Therefore, the simulation of a wind field has been performed where the wind turbines are modelled using an actuator disk. As in the shear-layer roll-up, the obtained $\Delta t$ with AlgEigCD for the FOM is consistently smaller than its counterpart for the ROM, proving the applicability of RedEigCD in the presence of time-dependent boundary conditions. In this case, the new self-adaptive timestepping method showcases a even larger improvement in timestep size compared to its FOM counterpart of up to a factor of 40 depending on the number of modes used and the flow conditions.
In both test cases, the obtained eigenbounds with the new method have been compared against the exact eigenbounds of the reduced operators as well as the estimated eigenbounds obtained by applying Gershgorin’s circle theorem to the reduced operators. The method showed good agreement with the exact eigenbounds and outperformed Gershgorin’s circle theorem estimates in both the shear-layer roll-up case and the actuator disk case. These results suggest that the method can provide accurate estimates of the eigenbounds of the reduced operators with and without non-homogeneous boundary conditions.
The present work has introduced RedEigCD, a stability-driven, self-adaptive timestepping framework that extends classical CFL and eigenvalue-based methods to reduced-order models of incompressible flow. Using the linear stability bounds derived from Poincaré’s separation theorem, it has been shown both theoretically and numerically that the stable timestep achievable in ROMs is systematically larger than their FOM counterparts. This theoretical stability guarantee, however, relies on the convective fields of the ROM and FOM being identical. In practice, the ROM trajectory naturally deviates from the FOM. Nonetheless, the stability of the FOM is determined by its smallest scales (large eigenvalues), which correspond to the modes truncated by the projection. Therefore, our proof justifies this increased stability in the reduced set of equations.
Beyond its theoretical insight, RedEigCD maintains the computational scalability of reduced models by confining all eigenvalue computations to the offline stage and by avoiding any FOM-size operations during the online stage. This results in an algorithm with a complexity of $\mathcal{O}(M)$, thus becoming negligible as the complexity of advancing the ROM in time is $\mathcal{O}(M^{3})$. The resulting algorithm thus provides a practical and generalizable timestep control mechanism for projection-based ROMs.
Future work will focus on the generalization of our method to the case where hyperreduction is employed for the convective terms, which would allow to reduce the online computational cost of the ROMs from $\mathcal{O}(M^{3})$ to $\mathcal{O}(M^{2})$. The main challenge will be to estimate the eigenbounds of the hyperreduced convective operator (the diffusive term remains unchanged). The practical performance of the method in more complex geometries and flow conditions should also be evaluated, but conceptually the method should be able to handle such situations. as it is general and can be applied to any flow configuration. Finally, the proposed RedEigCD framework can be extended to general non-quadratic problems by employing lifting transformations to recast the governing equations into a quadratic-bilinear form [54]. This approach eliminates non-quadratic nonlinearities, thereby restoring the algebraic structure to directly apply RedEigCD.

### v2

@@ 4 Numerical experiments
In this section, the new adaptive timestepping method outlined in Algorithm 1 is tested for two different fluid flow test cases. First of all, the robustness of the method in a case with periodic boundary conditions, namely the shear-layer roll-up, is presented. Moreover, the evolution of the maximum stable $\Delta t$ for different numbers of modes is presented, and compared with the maximum stable $\Delta t$ using the AlgEigCD. In the second case, the influence of non-homogeneous boundary conditions on the preservation of stability is considered for the unsteady wake behind an actuator disk [8]. In both test cases, the classical explicit fourth-order RK scheme is used in a two-dimensional incompressible Navier-Stokes solver [59, 60]. Both methods have been run for $M=[16,32,64,128,200]$ modes. Table 2 summarizes the methods used in this work.
@@
C ROM+constant ROM with constant $\Delta t$: $0.1$ for the shear-layer roll-up, $4\pi/200$ for the actuator disk [8]; used to compare accuracy of ROM+RedEigCD.
D Best ROM Best approximation ROM (Eq. (39)) with constant $\Delta t$; used to verify $\Delta t_{\text{ROM}}>\Delta t_{\text{FOM}}$ throughout simulation.
@@ 4.1 Shear-layer roll-up
Figure 6 (left) shows that the new self-adaptive timestepping method is effective in providing a larger timestep for the ROM than for the FOM, as expected from Theorem 1, as even for $M=200$, the largest number of modes considered, the ratio is around 1, which is indicated in the plot with a horizontal thicker line. Depending on the current flow physics and number of modes of the ROM, this timestep can be up to 9 times larger than for the FOM, which significantly accelerates the simulation by notably reducing the number of timesteps required to reach the final time. Furthermore, Figure 6 (left) shows how the ratio between the timesteps obtained in Case B and Case A converges to 1 for increasing ROM dimension, as expected from Theorem 1, as the eigenvalues of the reduced operators interlace with those of the full-order operators. Physically, this trend is driven by the inclusion of higher-index POD modes. As these capture the steeper spatial gradients of the flow, the ROM’s spectral radius monotonically approaches the FOM limit, as established in Section 3.2. This is especially clear for $M=200$, where the singular values have decayed to $\mathcal{O}(10^{-9})$, and the truncation errors between the ROM and the FOM (Figure 5) get closer to zero. Furthermore, note that at certain points, especially at the beginning of the simulation, the timestep ratio for $M=200$ is slightly smaller than 1.
@@
Increasing the number of modes, the timestep for the ROM is becoming closer to that of the FOM, as observed in Figure 6 (left) and 7. By doing so, steeper gradients, which are the ones driving the magnitude of eigenbounds, are captured by the operators. This indeed pushes the magnitude of the reduced eigenbounds towards the magnitude of the FOM eigenbounds, thus justifying this observed trend.
Figure 8 shows that the errors for both Case B and Case C have the same magnitude regardless of the number of modes ($M=16$, left; $M=200$, right). Thus, the use of the self-adaptive timestepping methods as performed in these cases does not significantly affect the accuracy of the ROM compared to the error of the best approximation ROM, as shown in Figure 8. In this case, $\Delta t$ for Case C is set to 0.01, as in the test case from [8]. To isolate the temporal accuracy of the stability-controlled ROM from the spatial truncation error, we define the reduced error relative to the projected FOM snapshots, $\mathbf{a}_{\mathrm{best}}(t)$ which is defined as
@@
where $\mathbf{a}_{\mathrm{test}}$ indicates the reduced velocity field used to compute the error, i.e., $\mathbf{a}_{\mathrm{\texttt{B}}}$, $\mathbf{a}_{\mathrm{\texttt{C}}}$. Figure 8 shows that either for $M=16$ or for $M=200$ there is no loss of accuracy for the new self-adaptive timestepper (Case B) compared to Case C, which indicates that in this test case there is no loss of accuracy regardless of the number of modes. However, in Case B the use of RedEigCD provides potential gains of up to a factor of 9 compared to a constant $\Delta t$.
Figure 8: Error comparison of the ROM results for $M=16$ (left) and $M=200$ (right) in the shear-layer roll-up of $\text{Re}=1000$ for Case B and C.
In terms of efficiency, the ROM with $M=64$ reduced the number of unknowns from $3\times 10^{4}$ to 64, while simultaneously allowing timesteps up to almost 5 times larger than the FOM. This reduced the total number of timesteps, while the wall-clock time for each timestep was also reduced due to dimensionality reduction, thus having benefits from both factors.
@@ 4.2 Flow through an actuator disk
In this case, the air flow around a wind turbine is modelled as in [23]. The simulation geometrical domain is $0,10]\times[-2,2]$, with a time-dependent prescribed inlet velocity at the left boundary, defined as follows,
@@
The other boundaries correspond to an outflow condition [59, Section 6.4.2]. The wind turbine has been modelled using an actuator disk of diameter 1, placed at $(2,0)$, which corresponds to a momentum sink modelled as a constant force $f=0.25$ acting in the negative $x$-direction. The initial conditions for this case are
@@
The magnitude of the ROM error, calculated with Eq.(40), is not affected by the variable $\Delta t$ compared to the reference result from [8], where $\Delta t$ was set to $4\pi/200$. As shown in Figure 11, the accuracy of the ROM+RedEigCD (Case B) is comparable to that of the ROM with constant $\Delta t$ (Case C), for both $M=16$ and $M=200$.
Figure 11: Error comparison of the ROM results for $M=16$ (left) and $M=200$ (right) in an actuator disk simulation with $\text{Re}=100$ for Case B and Case C.
@@ 4.3 Accuracy of the eigenbound estimates
To compute an overall error between the actual eigenvalues and the estimations, the following error is defined:
@@
4.4 Computational performance of RedEigCD
Assessing the performance of RedEigCD in terms of its wall-clock time requires analyzing the online computational efficiency as well as the offline overhead. While a complexity analysis on the costs is already performed and defined in Table 1, an assessment on actual runs is also provided here.
The online computational efficiency of the RedEigCD algorithm is evaluated by comparing the total online wall-clock time of the ROM against the FOM. We define the total speed-up factor, $\mathcal{S}$, as the product of the reduction in cost per iteration and the increase in the average timestep size:
$\mathcal{S}=\frac{T_{\mathrm{FOM}}}{T_{\mathrm{ROM}}}\frac{{\Delta t}^{\mathrm{ROM}}}{\Delta t^{\mathrm{FOM}}},$ (44)
where $T_{\mathrm{ROM}}$ and $T_{\mathrm{FOM}}$ denote the wall-clock time per iteration of the ROM and the FOM, respectively. Figure 13 visualizes these components for the shear-layer (left) and the actuator disk (right). The results indicate that while for small values of $M$ the computational complexity of $\mathcal{O}(M^{3})$ is acceptable, for the larger values tested, hyper-reduction would be a preferable option. However, RedEigCD shows benefits in the whole range of investigated modes.
Figure 13: Total speed-up of the RedEigCD algorithm for the roll-up of a shear-layer at Re=1000 (left) and the numerical simulation of a wind field of Re=100 (right).
On the other hand, regarding the offline costs, we define a cost metric $\mathcal{C}$ as the equivalent of ROM timesteps required to perform the offline setup of RedEigCD:
$\mathcal{C}=\frac{T_{\mathrm{offline}}}{T_{\mathrm{ROM}}},$ (45)
where $T_{\mathrm{offline}}$ corresponds to the total wall-clock time spent in the offline stage. For the most expensive case ($M=200$), this yields an offline cost of approximately 20 ROM iterations. While for extremely short simulations this overhead may be perceptible, for production runs involving thousands of timesteps, the cost of the offline stage becomes less than 1% of the whole simulation, which makes it insignificant.
@@ 5 Conclusion
In this paper, a novel self-adaptive timestepping framework, RedEigCD, has been presented for the time integration of POD-Galerkin ROMs of the incompressible Navier-Stokes equations. The methodology builds upon the conceptual foundations of EigenCD [6] and AlgEigCD [7], wherein the eigenbounds of the convective and diffusive operators are estimated to ensure the system remains at the edge of the stability region of the explicit time integration scheme. To preserve the online efficiency inherent to reduced-order modeling, RedEigCD avoids operations scaling with the FOM dimension by utilizing the sub-additivity of eigenbounds and Bendixson’s rectangle. By extending the decomposition approach of Sanderse [8] and Rosenberger and Sanderse [23], the reduced convective operator is split into quadratic and linear contributions, allowing the stability limit to be computed online as a simple linear combination of offline-calculated bounds.
Our proposed method specifically addresses the stability constraint, which defines the maximum allowable timestep to prevent numerical divergence. While in certain scenarios, particularly with implicit solvers, the maximum stable timestep can exceed that required for temporal accuracy, RedEigCD is tailored to explicit integration. Nonetheless, in cases where the accuracy limit may be exceeded, RedEigCD can be used in a dual-criterion approach.
The presence of non-homogeneous boundary conditions was addressed by splitting their contribution into symmetric and skew-symmetric parts. This allows the method to remain robust even when boundary interactions break the inherent skew-symmetry of the underlying symmetry-preserving ROM.
Numerical validation on a periodic shear-layer roll-up demonstrated that the $\Delta t$ obtained via RedEigCD is systematically larger than the FOM limit. This is theoretically supported by the interlacing of eigenvalues, which ensures smaller eigenbounds for the reduced operators. The benefits in timestep size for the ROM+RedEigCD for the shear-layer roll-up are up to a factor of 9, depending on the modal truncation and flow conditions, yielding significant computational speedups.
Furthermore, the framework was tested in a complex wind field simulation featuring time-dependent boundary conditions and actuator disks. The results confirmed that the ROM stability limit remains consistently larger than its FOM counterpart even under transient forcing. In this configuration, RedEigCD achieved improvements in timestep size of up to a factor of 40. In both test cases, the method demonstrated superior accuracy compared to Gershgorin’s circle theorem, providing a sharp and reliable estimate of the exact reduced eigenbounds.
Higher Reynolds numbers than those tested increase the spectral radius and tighten stability limits. In this situation, we expect RedEigCD to remain robust as numerical stability is governed by the eigenvalues of the linearized system. Though nonlinear stability analysis would offer a global perspective, the linear spectral bounding used here provides a more computationally efficient proxy for real-time adaptive timestepping. Flow regimes with high Reynolds number are often characterized by a slow decay of singular values, requiring a larger number of modes $M$ that severely increases the computational cost. Even though our method, which has been tested for cases with a faster decay in this study, will still provide the maximum stable timestep, the benefits in terms of timestep will be smaller in cases that require a larger number of modes, leading to timesteps closer to those of the FOM.
A key theoretical takeaway of this work is the formal connection established between linear stability theory and reduced-order modeling. Based on the combined theorems of Bendixson [39] and Rao [40], it has been proven that the stable timestep achievable in a projection-based ROM is always larger than or equal to that of the FOM. This is justified by the fact that FOM stability is dictated by the smallest spatial scales (associated with the largest eigenvalues), which are precisely the components truncated during the POD projection. RedEigCD preserves ROM scalability with an online complexity of $\mathcal{O}(M)$, making the cost of the adaptive controller negligible compared to the $\mathcal{O}(M^{3})$ cost of the state advancement.
Future work will explore the generalization of this framework to hyperreduced convective terms to further reduce online costs to $O(M^{2})$. Additionally, the method will be evaluated in more complex geometries to test its robustness in highly irregular spectral environments. Finally, the proposed RedEigCD framework can be extended to a wide range of non-quadratic systems that admit a quadratic-bilinear representation through lifting transformations [61], which recast many smooth nonlinearities into a form where RedEigCD can be directly applied.